# Excitation Amplitude Sampling for Low Variance Electronic Structure on Quantum Computers

### v1

@@ Appendix D Direct configurational sampling
In Fig. 7(b) we also show the standard deviation for these different energy estimates using 1000 shots at different system sizes with a fixed interatomic distance of $1.5$Å. The combination of an additional polynomial factor that arises from needing to split the shot budget between an increasing number of observables and the increased dependence on $c_{0}$ arising from calculating $c_{0}c_{ij}^{ab}$ factors leads to a substantially worse scaling with system size compared to the classical shadow approach in almost all scenarios. This validates the benefits found from the combination of shadow sampling and the mixed estimator presented in the main text.

### v2

@@ Appendix E Direct configurational sampling
In Fig. 8(b) we also show the standard deviation for these different energy estimates using 1000 shots at different system sizes with a fixed interatomic distance of 1.51.5Å. The combination of an additional polynomial factor that arises from needing to split the shot budget between an increasing number of observables and the increased dependence on $c_{0}$ arising from calculating $c_{0}c_{ij}^{ab}$ factors leads to a substantially worse scaling with system size compared to the classical shadow approach in almost all scenarios. This validates the benefits found from the combination of shadow sampling and the mixed estimator presented in the main text.